# L4Q: Parameter Efficient Quantization-Aware Fine-Tuning on Large Language Models

### v4

@@ Appendix F Training Time
QA-LoRA 5.0 11.2 19.8 36.2
L4Q 4.4 10.1 16.9 39.6
Appendix G Throughput and Speedup of fully-quantized models and mixed-precision models
@@
Appendix H Detailed Result on Main Evaluations
We present the Commonsense QA and MMLU benchmark results with averaged accuracy score on Section 4. We present the detailed results of each benchmarks composed of several categories of tasks in Table 11 and Table 12 below. Through evaluation, we demonstrate that L4Q generally achieves higher accuracy in low-bit quantized models compared to both PTQ methods and PTQ-based fine-tuning methods. Notably, L4Q surpasses the pre-trained models on the Commonsense QA benchmarks and on the MMLU benchmarks with LLaMA-1 7B and 33B models. In contrast, PTQ-based fine-tuning methods, including those that incorporate high-precision LoRA weights, show lower performance compared to both L4Q and the pre-trained models. These results emphasize the challenges of recovering from quantization errors with PTQ alone and highlight the effectiveness of L4Q’s joint quantization and fine-tuning scheme.
Table 11: Commonsense QA benchmark result. The numbers represent accuracy (%) of each task.
@@ Appendix H Detailed Result on Main Evaluations
LLaMA 33B None 16 63.3 81.0 52.8 80.4 75.9 82.6 36.0 67.4
@@
Table 12: MMLU benchmark result. The numbers represent accuracy(%) of each task.
@@
LLaMA 2 13B None 16 47.8 42.3 60.5 59.4 52.1 52.0 43.8 63.0 61.2 54.8

### v5

@@ E.2 Training Time
QA-LoRA 5.0 11.2 19.8 39.6
L4Q 4.4 10.1 16.9 37.9
Appendix F Throughput and Speedup of fully-quantized models and mixed-precision models
@@
Appendix G Detailed Result on Main Evaluations
We present the Commonsense QA and MMLU benchmark results with averaged accuracy score on Section 4. We present the detailed results of each benchmarks composed of several categories of tasks in Table 13a to Table 14b below. Through evaluation, we demonstrate that L4Q generally achieves higher accuracy in low-bit quantized models compared to both PTQ methods and PTQ-based fine-tuning methods. Notably, L4Q surpasses the pre-trained models on the Commonsense QA benchmarks and on the MMLU benchmarks with LLaMA-1 7B and 33B models. In contrast, PTQ-based fine-tuning methods, including those that incorporate high-precision LoRA weights, show lower performance compared to both L4Q and the pre-trained models. These results emphasize the challenges of recovering from quantization errors with PTQ alone and highlight the effectiveness of L4Q’s joint quantization and fine-tuning scheme.
Table 13a: Commonsense QA benchmark result. The numbers represent accuracy (%) of each task.
@@ Appendix G Detailed Result on Main Evaluations
Table 13b: Commonsense QA benchmark result. The numbers represent accuracy (%) of each task.
Model Method #Bits Hella. PIQA ARC-c ARC-e Winogr. BoolQ OBQA Avg.
LLaMA-1 33B None 16 63.3 81.0 52.8 80.4 75.9 82.6 36.0 67.4
@@
Mistral-v0.1 7B None 16 61.2 80.6 50.4 80.9 73.9 83.6 32.6 66.2
LoRA 16 61.2 82.1 50.3 80.9 74.0 83.7 32.6 66.4
GPTQ 4 59.8 82.3 46.9 79.4 73.5 84.4 30.6 65.3
OmniQ 4 60.7 79.9 47.1 78.2 73.6 82.5 31.2 64.7
LoftQ* 4&16 54.2 78.0 44.5 77.6 67.5 75.1 27.8 60.7
QLoRA* 4&16 60.8 82.1 50.5 80.4 73.2 81.8 32.0 65.8
QA-LoRA 4 60.6 81.7 49.0 79.5 73.3 81.8 32.0 65.4
QAT-LoRA 4&16 60.3 80.0 46.8 78.6 73.4 82.6 29.8 64.5
L4Q 4 60.3 81.6 50.9 80.4 72.4 84.8 32.0 66.1
GPTQ 3 57.3 79.5 43.5 75.8 70.6 78.0 27.8 61.8
OmniQ 3 58.7 79.1 43.4 75.2 69.8 72.3 31.0 61.4
LoftQ* 3&16 55.2 74.7 40.9 72.4 63.8 76.5 26.2 58.5
QLoRA* 3&16 57.5 80.3 46.6 76.7 69.8 80.7 29.6 63.0
QA-LoRA 3 57.4 78.7 43.9 75.7 70.9 80.9 28.4 62.3
QAT-LoRA 3&16 56.8 79.7 40.9 74.6 70.3 79.5 29.4 61.6
L4Q 3 57.5 80.2 47.7 78.0 66.5 83.8 28.4 63.1
Table 14a: MMLU benchmark result. The numbers represent accuracy(%) of each task.
@@
Table 14b: MMLU benchmark result. The numbers represent accuracy(%) of each task.
0-shot 5-shot
LLaMA-2 13B None 16 47.8 42.3 60.5 59.4 52.1 52.0 43.8 63.0 61.2 54.8
@@
Mistral-v0.1 7B None 16 54.1 51.2 70.5 67.8 60.2 56.5 52.6 73.5 70.4 62.6
LoRA 16 54.5 51.4 70.9 68.2 60.6 56.8 52.9 73.9 70.8 62.9
GPTQ 4 51.8 47.4 68.1 65.5 57.6 55.9 50.4 71.6 68.1 61.0
OmniQ 4 52.4 49.2 69.0 65.7 58.4 55.1 51.8 71.2 68.5 61.0
LoftQ* 4&16 38.4 40.8 53.5 51.1 45.2 39.9 41.6 53.4 50.8 45.7
QLoRA* 4&16 52.6 49.2 69.6 66.0 58.7 55.7 51.7 71.2 68.0 61.1
QA-LoRA 4 51.4 46.4 66.0 64.1 56.5 56.1 51.4 72.0 67.6 61.2
QAT-LoRA 4&16 52.1 48.8 69.5 67.7 58.8 54.1 50.5 71.0 68.0 60.2
L4Q 4 52.6 48.9 69.9 67.1 59.0 56.3 51.1 71.7 68.8 61.4
GPTQ 3 45.0 42.4 59.7 57.2 50.5 43.0 43.0 57.4 57.8 49.6
OmniQ 3 49.4 44.9 63.4 61.6 54.3 50.0 47.0 66.1 63.0 55.9
LoftQ* 3&16 32.2 33.3 38.0 41.3 35.8 34.6 34.1 39.6 41.1 37.1
QLoRA* 3&16 46.9 43.4 60.6 60.2 52.2 47.6 44.8 61.6 62.9 53.6
QA-LoRA 3 45.8 42.9 60.8 54.5 50.5 48.8 43.6 59.2 56.4 51.7
QAT-LoRA 3&16 47.6 43.0 61.3 59.9 52.4 48.5 46.3 63.6 60.2 54.0
L4Q 3 49.9 43.8 63.0 63.0 54.5 51.1 46.1 66.4 62.5 56.2